OpenStudy (jojo12g):

Brahim completed 18 free throws and missed 12. Using experimental probability, how many free throws is he predicted to complete in the next 100 tries? A. 6 B. 36 C. 48 D. 60

OpenStudy (jojo12g):

i will give medal and fan

OpenStudy (omarbirjas):

This one is easy. All we have to do is setup a ratio and cross multiply.

OpenStudy (omarbirjas):

What is the sum of 18 and 12?

OpenStudy (anonymous):

\[\frac{18}{18+12}=\frac{x}{100} \]x = 60